# Coherently Coupled Carrier and Phonon Dynamics in Elemental Tellurium Probed by Few-Femtosecond Core-Level Transient Absorption

### v1

@@ V Summary
Using few-femtosecond core-level transient absorption at the Te N4,5 edge we have observed a coherent coupling of the photo-excited holes and electrons to the A1g phonon mode in elemental tellurium. Following excitation of the lone-pair orbitals to the anti-bonding conduction band, the lattice is excited along the A1g phonon coordinate to a new metastable excited state that lives on the timescale of carrier recombination. This excited lattice configuration is found to follow the 31(2) ps carrier population decay, while hot carrier relaxation () driven by electron-phonon coupling is much faster and occurs over 1.60(2) ps. Coherent excitation of the A1g phonon mode modulates the hot carrier temperature (), in which when the phonon displacement is maximized (Te symmetrized), the carrier temperature is minimized due to an increase in electronic heat capacity as Te becomes more metallic and the density of states is increased at the Fermi level. This back-coupling of the lattice energy to the carriers is observed for the first time in a semiconductor and is driven by a coupling between the electronic and lattice degrees of freedom along the A1g coordinate. The results are corroborating evidence of the light-induced semiconductor-to-metal phase transition [8, 10, 12] by demonstrating a strong dependence of the electronic density of states on . While depopulating the lone-pairs and populating the anti-bonding conduction band leads to the metastable Te phase, Fourier analysis demonstrates the launching of the coherent phonon is not through a pure displacive excitation, but has a weak dependence upon carrier relaxation possibly due to energy-dependent electron-phonon coupling. The measurement of coherent coupling between the hot carriers and A1g phonon mode in elemental Te provides the basis for further investigation of the light-induced phase transition from a semiconductor to metal and how it may couple to other degrees of freedom, such as spin or the non-totally symmetric phonon modes.
@@ VI Acknowledgments
Appendix A Experimental Appartus
The XUV transient absorption experiments are conducted on a table-top apparatus that starts with a Ti:sapphire laser (Coherent Astrella). The output is 7 mJ at 1 kHz with a pulse duration of 35 fs at a center wavelength of 795 nm. From the laser output, 2.3 mJ is reflected from a beam splitter and focused into a 2.2 meter long 400 micron inner diameter hollow core fiber filled with neon at a pressure of 30 Psi for supercontinuum generation. The supercontinuum output of the hollow core fiber spans an octave from 500 to 1000 nm with a pulse energy of 1.5 mJ and is mechanically chopped to 100 Hz prior to collimation. The beam is subsequently collimated with a magnification of 1.25 to the fiber input and compressed with seven pairs of PC70 double angle chirp mirrors (Ultrafast Innovations) that support the entire beam bandwidth and a Z-cut 2 mm thick ammonium dihydrogen phosphate crystal [47]. The resulting beam is sent into the Mach-Zender interferometer via a 1 mm thick beam splitter (90T:10R), where the transmitted beam is used for HHG, and the reflected portion is used for the pump. Both pump and HHG driving pulse are optimized with fused silica wedges for fine tuning of the dispersion and are approximately 4 fs in duration. The HHG driving arm is focused (50 cm focal length) into an argon filled gas cell for HHG where a broadband XUV (35-73 eV) pulse is generated as the probe pulse. The residual NIR-VIS driving field is filtered out with a 100 nm thick Al foil that acts as a high-pass filter. The XUV pulse is focused onto the thin film sample by a gold-coated toroidal mirror implemented in a 2f-2f imaging configuration and subsequently dispersed onto a CCD camera (PIXIS 400B) in a flat-field grating spectrograph where the transmitted spectrum is measured. The pump in the Mach-Zender is time delayed to the probe by an optical delay line with an optical encoder (Attocube), and focused (1 m focal length) onto the sample after recombining in a collinear geometry with the XUV by an annular mirror. The pump beam is removed by a second 100 nm thick Al foil prior to entering the spectrograph chamber. The pump beam is both transmitted and blocked at each time delay to collect XUV spectrum with and without the pump. The XUV absorption spectra undergo noise reduction via an edge-referencing scheme [48] in the 60-73 eV region where no TA signal is observed to improve the noise floor to about 1 mOD.
@@ Appendix C Instrument response function
Appendix D Residual Transient Absorption
@@
The residual TA signal is presented in Fig. 9 where there is a few mOD signal near the 4d5/2 to Fermi level (40.4 eV) transition that rapidly decays on the timescale of a couple hundred femtoseconds. This residual increase in absorption is well below the Fermi level, and a decrease near the Fermi level is consistent with non-thermalized holes and electrons not captured by the SVD that we assign to the hot carriers. Additionally, the hot carriers dissipate their energy into low energy carriers near the Fermi level that persist to 31 ps , which results in the appearance of state blocking/filling contributions emerging in the residual TA after the 1.6 ps decay. In addition to evolution of the state blocking via the hot carriers, the phase variation in energy of the phonon motion is not possible to capture with a single SVD component and is thus found in the residual TA. This phase variation in energy of the A1g phonon motion results in oscillatory features in the residual TA, alongside potential features of the E1TO mode excitation.
Appendix E Low fluence measurement
@@
In Fig. 10 we present the transient absorption spectra at a carrier density of . At this lower carrier density, below the expected threshold for the photo-induced semiconductor-to-metal phase transition, the phonon induced edge shift is clearly observed, however, the substantially smaller state blocking/filling contributions are not resolved. The SVD applied to these measurements reveals a single SVD vector resembling the phonon induced absorption edge shift. All other SVD components are indistinguishable from noise. From this, we find no measurable decay () of the excited state equilibrium position and coherent phonon () on the timescale of the measurement. The retrieved frequency is 3.6(3) THz, closer to the equilibrium frequency due to a reduced electronic softening.
Appendix F Short timescales
As discussed above in section IV, the shift of the Te equilibrium position in the DECP model should depend on the carrier population [15]. In Fig. 11 we show the phonon induced edge shift component of the multi-temperature model () to a measurement utilizing 660 as time-steps, with the longtime scale components fixed to the results retrieved from the scan to longer timescales. In conjunction with the Te measurement, TA of the neon Rydberg resonances are done after each scan cycle to correct for zero-delay drifts and this enables coupled fitting to retrieve dynamics on the timescale of the cross-correlation. Applying the same SVD methodology as above, we do a simultaneous fit of the Te edge shift and neon Rydberg resonances, where the instrument response convolution is a shared parameter. Through this, we retrieve a 3(1) fs rise time () of the edge shift describing the initial equilibrium position displacement of Te. This would suggest the response of is not immediate, but there is a time dependence of the displacive force. The tail of the signal is not captured by the fitting, which is attributable to a weak pre-pulse in the pump beam, visible at around 10 fs in Fig. 8(a).

### v2

@@ V Summary
Using few-femtosecond core-level transient absorption at the Te N4,5 edge we have observed a coherent coupling of the photo-excited holes and electrons to the A1 phonon mode in elemental tellurium. Following excitation of the lone-pair orbitals to the anti-bonding conduction band, the lattice is excited along the A1 phonon coordinate to a new excited state that lives on the timescale of carrier recombination. This excited lattice configuration is found to follow the 33(5) ps carrier population decay, while hot carrier relaxation () driven by electron-phonon coupling is much faster and occurs over 1.59(3) ps. Coherent excitation of the A1 phonon mode modulates the hot carrier temperature (), in which when the phonon displacement is maximized (Te more symmetric), the carrier temperature is minimized due to an increase in electronic heat capacity as Te becomes more metallic and the DOS is increased at the Fermi level. This back-coupling of the lattice energy to the carriers is observed for the first time in a semiconductor and is driven by a coupling between the electronic and lattice degrees of freedom along the A1 coordinate. These results that demonstrate a dependence of the electronic DOS on are consistent with a coherent phonon driven tunability of the electronic structure [8, 10, 12]. While depopulating the lone-pairs and populating the anti-bonding conduction band leads to the excited state Te structure, Fourier analysis demonstrates the launching of the coherent phonon is not through a pure displacive excitation, but has a weak dependence upon carrier relaxation possibly due to energy-dependent electron-phonon coupling. The measurement of coherent coupling between the hot carriers and A1 phonon mode in elemental Te provides the basis for further investigation of the light-induced increase in metallicity and how it may couple to other degrees of freedom, such as spin or the non-totally symmetric phonon modes.
@@ VI Acknowledgments
Appendix A Experimental Apparatus and Data Processing
The XUV transient absorption experiments are conducted on a table-top apparatus that starts with a Ti:sapphire laser (Coherent Astrella). The output is 7 mJ at 1 kHz with a pulse duration of 35 fs at a center wavelength of 795 nm. From the laser output, 2.3 mJ is reflected from a beam splitter and focused into a 2.2 meter long 400 micron inner diameter hollow core fiber filled with neon at a pressure of 30 Psi for supercontinuum generation. The supercontinuum output of the hollow core fiber spans an octave from 500 to 1000 nm with a pulse energy of 1.5 mJ and is mechanically chopped to 100 Hz prior to collimation. The beam is subsequently collimated with a magnification of 1.25 to the fiber input and compressed with seven pairs of PC70 double angle chirp mirrors (Ultrafast Innovations) that support the entire beam bandwidth and a Z-cut 2 mm thick ammonium dihydrogen phosphate crystal [52]. The resulting beam is sent into the Mach-Zender interferometer via a 1 mm thick beam splitter (90T:10R), where the transmitted beam is used for HHG, and the reflected portion is used for the pump. Both pump and HHG driving pulse are optimized with fused silica wedges for fine tuning of the dispersion and are approximately 4 fs in duration. The HHG driving arm is focused (50 cm focal length) into an krypton filled gas cell for HHG where a broadband XUV (35-73 eV) pulse is generated as the probe pulse. The residual NIR-VIS driving field is filtered out with a 100 nm thick Al foil that acts as a high-pass filter. The XUV pulse is focused onto the thin film sample by a gold-coated toroidal mirror implemented in a 2f-2f imaging configuration and subsequently dispersed onto a CCD camera (PIXIS 400B) in a flat-field grating spectrograph where the transmitted spectrum is measured. The pump in the Mach-Zender is time delayed to the probe by an optical delay line with an optical encoder (Attocube), and focused (1 m focal length) onto the sample after recombining in a collinear geometry with the XUV by an annular mirror. The pump beam is removed by a second 100 nm thick Al foil prior to entering the spectrograph chamber. The pump beam is both transmitted and blocked at each time delay to collect XUV spectrum with and without the pump.
The TA spectra are calculated at each time delay as , where and are the transmitted XUV with the NIR pump on and off, respectively. The XUV TA spectra are subject to de-noising via an edge-referencing scheme where no TA signal is observed to improve the noise floor to about 1 mOD [53]. This is done for a total of 93 cycles, over 465 time delays. Included in Fig. 5(a) is the de-noised TA without subtraction of the pre-zero delay heat background. To subtract the time-independent heat background, the first five negative time delays are averaged in time and subtracted from the TA spectra. The raw unprocessed data is shown in Fig. 5(b), where there is prominent high frequency noise.
@@ Appendix C Instrument response function
Appendix D Analytical Decomposition
We analyze here using an analytical model predominant spectral features in the transient absorption spectra at 150 fs, corresponding to the first oscillation maximum of the phonon. Well demonstrated in previous core-level TA measurements of elemental semiconductors [30] and transition metal dichacolgenides [27, 56], the excited state can be decomposed into individual contributions:
incorporating the phonon and band-gap renormalization induced absorption edge shift (), excited state inhomogeneous broadening of the core-level transition (), and carrier induced state blocking (filling) in the conduction (valence) band (), also called state removal/state opening. In the region of the Te N4,5 edge we show the decomposition into these components in Fig. 9. To retrieve the individual contributions, an optimization is performed at 1 eV around the Fermi level and the fitting errors retrieved via the Hessian matrix.
In this analytical model the absorption edge shift is optimized to a 580(30) meV uniform energy shift (rigid shift approximation) with respect to the ground state XUV absorption. An optimized 8.3(4) % scaling factor is applied to the ground state XUV absorption to obtain agreement with the A. We believe the scaling factor is due to non-uniform shifting and restructuring of the conduction band DOS across the BZ following photoexcitation [30, 18, 31]. The shift includes both changes to the conduction band energy and 4d level binding energies that we cannot independently discern. The inhomogeneous broadening of the excited state is achieved by convolving the static absorption with a Gaussian and yields a broadening meV. As hot carrier thermalization is expected to occur mostly within the first 100 fs, we approximate the carrier induced state filling and blocking as a Fermi-Dirac distribution over the DOS [24] with 0.5 eV lifetime broadening. This is replicated for the spin-orbit split core-levels where there is strong spectral overlap between the holes of the and electrons of the . From this, we estimate a shared electronic temperature of 3300(300) K for both holes and electrons at the maximum of the phonon displacement. Due to spectral congestion at the electron state blocking signal, the assumption is made of a shared electronic temperature. Additionally, the quasi-Fermi levels of the carriers are offset by half the energy of a pump photon (0.8 eV) from the ground state Fermi level [30]. These approximations are made to reduce the over parametrization of the optimization. However, this analytical model slightly over parametrizes the data, with a , but has a benefit of isolating the main spectral features.
While reproducing the change in absorption near the Fermi level, the rigid shift approximation made in the model causes deviations from experiment due to an overestimate of the edge shift at high energies (above 42 eV). This we attribute to a non-uniform electron-phonon coupling over a large photon energy range, as not all bands renormalize equally [30, 18]. In addition to the rigid shift approximation, the use of a scaling factor in the model warrants further consideration. When considering how the conduction band DOS responds to the A1 phonon induced lattice distortion, it is not a rigid red shift of all conduction band DOS, but the appearance of a small component towards the Fermi level while depleting the near CBM DOS. There is also considerable restructuring over the entire DOS [31]. This is in contrast to changes to the DOS in Bi upon a 1% increase in inter-atomic distances [18], where there is limited restructuring and it resembles a linear red shift of the conduction band in the vicinity of the Fermi level. In the case of Te, particularly under large lattice distortions, it is thus no surprise that the rigid shift approximation in the absence of a scaling factor breaks down. The scaling factor would suggest only a fraction of the DOS undergoes shifting in response to photoexcitation, consistent with how the DOS evolves with large lattice distortions [31]. The 4d core-level potential should, however, follow the traditional methodology of a rigid shift. Future investigations might consider a more advanced methodology, dependent upon detailed TDDFT calculations of how the DOS evolve with increasing lattice distortion, local field effects [34], and the core-level potential slope, and their changes can be approached in an independent manner.
Although this approach provides precise quantities within the approximations, lack of detailed knowledge of complex band dynamics and strong spectral overlap of lattice and carrier features gives rise to significant uncertainty. Further the necessity to utilize a scaling factor in the rigid shift approximation due to an unknown dependence of the shift on band dispersion, energy, and DOS restructuring impedes successful interpretations with this decomposition method.
Appendix E Residual Transient Absorption
@@
The residual TA signal is presented in Fig. 10 where there is a few mOD signal near the 4d5/2 to Fermi level (40.4 eV) transition that rapidly decays on the timescale of a couple hundred femtoseconds. This residual increase in absorption is well below the Fermi level, and a decrease near the Fermi level is consistent with non-thermalized holes and electrons not captured by the SVD that we assign to the hot carriers. While two-photon absorption could occur producing non-thermal holes, no increase in absorption discernible from the noise level is observed below below 38.5 eV, which exceeds the single photon energy of the NIR pump minus the bandgap. Additionally, the hot carriers dissipate their energy into low energy carriers near the Fermi level that persist to 33 ps, which results in the appearance of state blocking/filling contributions emerging in the residual TA after the 1.6 ps decay. In addition to evolution of the state blocking via the hot carriers, the phase variation in energy of the phonon motion is not possible to capture with a single SVD component and is thus found in the residual TA. This phase variation in energy of the A1 phonon motion results in oscillatory features in the residual TA.
Appendix F Low fluence measurement
@@
In Fig. 11 we present the transient absorption spectra at a carrier density of . At this lower carrier density, below the expected threshold for the photo-induced semiconductor-to-metal phase transition, the phonon induced edge shift is clearly observed, however, the substantially smaller state blocking/filling contributions are not resolved. The SVD applied to these measurements reveals a single SVD vector resembling the phonon and electronic screening induced absorption edge shift. All other SVD components are indistinguishable from noise. From this, we find no measurable decay () of the excited state equilibrium position and coherent phonon () on the timescale of the measurement. The retrieved frequency is 3.6(3) THz, closer to the equilibrium frequency due to a reduced electronic softening.
Appendix G Short timescales
As discussed above in section IV, there is an early timescale rise of the energy shift trace that is too fast to be explained by coherent phonon motion. In Fig. 12 we show the energy shift trace component of the SVD analysis to a measurement utilizing 660 as time-steps, and the TA signal integrated over the energy ranges 40.35-40.5 eV and 38.5-40 eV. In conjunction with the Te measurement, TA of the neon Rydberg resonances are done after each scan cycle to correct for zero-delay drifts, where the standard deviation of zero-delay drifts over 113 scan cycles is found to be 508 as, indicating minimal long-term drifts. The energy range of 40.35-40.5 eV can be mostly attributed to the absorption edge energy shift contribution based on the analytical model (see Appendix D), while the 38.5-40 eV range is the holes. We find the rapid rise of the energy shift vector is near instantaneous with the cross-correlation, suggesting a rapid electronic response that contributes to a red shift of the Te N4,5 absorption edge, followed by the start of a slow-rise due to the onset of phonon motion. This is consistent with the energy integrated TA signal over the 40.35-40.5 eV region. The energy integrated TA signal over the hole region (38.5-40 eV) exhibits a rise that follows the cross-correlation, but maintains a near constant integrated signal outside of the cross-correlation on these short timescales. Despite the pump field intensity being on the order of the approach to non-linear effects - in semiconductors such as Ge [57] and GaAs, [58] non-linear excitation has been observed via XUV - we do not find in the integrated TA signal over the hole region any dynamics associated with non-linear excitation effects in the measurements. This outcome could be due to attosecond timing instability during an individual scan cycle and there may be a need to sample at a higher frequency. Additionally, the weak pre-time-zero delay signal outside the cross-correlation is attributable to a weak pre-pulse in the pump beam, visible at around 10 fs in Fig. 8(a).